# Trial by FIRE: Probing the dark matter density profile of dwarf galaxies with GraphNPE

### v1

@@ 2.2 Test samples of dwarf galaxies
Given that our goal is to apply GraphNPE on Milky Way satellite galaxies, we will focus our studies to the satellites of the Milky Way-mass simulated galaxies. In FIRE-2, halos and subhalos are identified using a modified version of the rockstar  [^5] https://bitbucket.org/awetzel/rockstar-galaxies six-dimensional halo finder (Behroozi et al., 2013), which accounts for multi-mass and multi-species particles. For each subhalo, we use the DM particles assigned by rockstar and identify its associated star particles. We use a similar procedure to Necib et al. (2019) to assign member star particles: we require that the star particles lie within the current $r_{\mathrm{200m}}$ (the radius within which the average density is 200 times the mean matter density of the Universe) of the subhalo, and have velocities within $3\sigma$ of the subhalo’s stellar velocity dispersion
@@
In addition, we select subhalos with halo-to-stellar mass ratios satisfying $M_{\mathrm{halo}}/M_{\mathrm{\star}}>10^{2}$ and with the number of star particles between $N_{\mathrm{\star}}\in[20,5000]$. This roughly corresponds to a stellar mass range of approximately $M_{\mathrm{\star}}\in[10^{5},10^{7}]\,\mathrm{M_{\odot}}$ and a halo mass range of $M_{\mathrm{halo}}\in[10^{7},10^{10}]\,\mathrm{M_{\odot}}$. [^6] Peak halo mass range of $M^{\mathrm{peak}}_{\mathrm{halo}}\in[10^{8},10^{10}]\,\mathrm{M_{\odot}}$. The number of stars are chosen to encompass observations of classical and faint dwarfs. For instance, the highest number of resolved stars observed for Milky Way dwarfs, Fornax and Sculptor, are 2483 and 1365 stars, respectively (Walker et al., 2009a).
Next, we center each subhalo, and compute the positions and velocities of star and DM particles of each subhalo relative to the subhalo’s DM barycenter. We identify the center of each subhalo using the shrinking sphere method (Power et al., 2003) on the DM particles only. Starting with an initial guess for the center, we iteratively compute the DM barycenter of all particles within a specified radius. At each step, the radius is reduced by 10%, and the center is updated to the newly calculated barycenter. The iterations proceed until the center converges to within 0.1% of the previous iteration. In rare cases, when a subhalo is located too close to its host or another subhalo, the computed center may still be offset. These cases are manually checked. Additionally, the stellar and DM centers are not always perfectly aligned, potentially complicating the analysis. For simplicity, we assume these centers coincide.
@@ 3.1 Training Simulation & Forward Model
We assume the galaxies to be spherically symmetric and in dynamical equilibrium (i.e., the distribution function $f(\vec{x},\vec{v};\theta)$ does not explicitly depend on time). These are common assumptions in traditional Jeans analysis, and as previously discussed, might not be robust against realistic Milky Way dwarfs. Ideally, one can construct a more realistic family of distribution functions from hydrodynamic simulations, such as FIRE-2. However, such simulations are computationally demanding and currently impractical for generating the large training sets required for our framework. Moreover, as we will demonstrate in Section 7.1, even high-resolution zoom-in simulations struggle to capture the observed population of dwarf galaxies, particularly the ultra-faints.
@@
where $\sigma_{r}$ and $\sigma_{t}$ are the radial and tangential velocity dispersions. The velocity anisotropy ranges from $\infty$- ∞ to $1$, where $\beta=1,0,-\infty$ corresponds to a radial, isotropic, and tangential velocity profile, respectively. We assume a functional form for $\beta(r)$ following the Osipkov-Merritt (OM) profile (Osipkov, 1979; Merritt, 1985),
@@
where $\beta_{0}$ is the normalization and $r_{a}$ is the anisotropy scale radius. The scale radius $r_{a}$ thus determines the transition from $\beta_{0}$ at small radii to a radially-biased orbits at larger radii. In contrast to the N23 model, which always set $\beta_{0}$ to 00, we let $\beta_{0}$ be a free parameter within the range $\left[-0.5,1\right]$.
@@
$\log_{10}(r_{\mathrm{dm}}/\mathrm{kpc})$ DM scale radius gNFW (Eq. 2) $[-2,2]$
$\gamma$ DM inner slope gNFW (Eq. 2) $[-1,2]$
@@
$\beta_{0}$ Velocity anisotropy normalization OM (Eq. 6) $[-0.5,1]$
@@
In total, the current model has three DM parameters ($\rho_{0}$, $r_{s}$, $\gamma$) and three stellar parameters ($\beta_{0}$, $r_{a}$, $r_{\star}$). We expand the prior range for these parameters from N23 and summarize them in Table 2. Additionally, we select only galaxies with 3-D velocity dispersions within $0.1-50\,\mathrm{km/s}$ for the final training set, which should sufficiently include the population of classical and ultra-faint dwarfs (Pace, 2024).777 Assuming an isotropic velocity distribution, the 3-D velocity dispersion $\sigma_{\mathrm{3-D}}$ is related to the line-of-sight velocity dispersion $\sigma_{\mathrm{los}}$ via $\braket{\sigma_{\mathrm{3-D}}^{2}}=3\braket{\sigma_{\mathrm{los}}^{2}}$. For $\sigma_{\mathrm{3-D}}$ in the range of $(0.1-50)\,\mathrm{km/s}$, $\sigma_{\mathrm{los}}$ is approximately within $(0.06-28.9)\,\mathrm{km/s}$. Among Milky Way dwarfs, the highest and lowest observed line-of-sight velocity dispersions–excluding the Small and Large Magellanic Clouds (SMC and LMC)–is $12.1\pm 0.2\,\mathrm{km/s}$ for Fornax (Walker et al., 2009a; Górski et al., 2011; Muñoz et al., 2018; Wang et al., 2019) and $1.2^{+0.9}_{-0.6}\,\mathrm{km/s}$ for Tucana V (Hansen et al., 2024).
@@ 4.1.1 DM density and velocity anisotropy profiles
Figure 2: Comparison between the inferred velocity anisotropy profiles $\beta(r)=1-\sigma^{2}_{t}/\sigma^{2}_{r}$, where $\sigma_{t}$ and $\sigma_{r}$ are the tangential and radial velocity dispersions, from GraphNPE (orange) and Jeans modeling (blue) for a selection of FIRE-2 galaxies. Values of $\beta=0,1-\infty$ corresponds to isotropic, radially-biased, and tangentially-biased orbits. Panels are the same as in Figure 1.
Figure 1 and Figure 2 show the recovered DM density profiles $\rho_{\mathrm{dm}}(r)$ and velocity anisotropy profiles $\beta(r)$ for a subset of the selected galaxies, with the remaining profiles provided in Appendix C.1. Each panel shows a single dwarf galaxy, labeled with its rockstar halo ID at $z=0$ and a corresponding letter (e.g., A, B, C, D) for readability, and compares the profiles derived using GraphNPE (blue) and Jeans modeling (orange). The posterior median profiles are shown as solid lines, with shaded bands indicating the 68% and 95% confidence intervals, while black circles with error bars represent the true profiles. [^8] Note that error bars are not visible for the true density profiles. The dashed vertical lines indicate $(0.25,1,4)\,r_{\mathrm{1/2}}$, where $r_{\mathrm{1/2}}$ is the 3-D half-stellar mass radius.
@@
Figure 3: Comparison between the inferred mass and structural parameters from GraphNPE (orange) and Jeans modeling (blue). From left to right, the columns show the peak circular velocity $V_{\mathrm{max}}$, the half-stellar radius mass $M_{r_{\star,1/2}}$, and the virial mass $M_{\mathrm{200m}}$. $V_{\mathrm{max}}$ and $M_{\mathrm{200m}}$ are calculated by extrapolated the DM density profiles to $r_{\mathrm{max}}$ and $r_{\mathrm{200m}}$, respectively. The top and bottom rows show the median and fractional residual of the predicted values against the true values, respectively, with the error bars denoting the 64th percentile ranges. The black dashed line represents the one-to-one correlation. Since the virial mass $M_{\mathrm{200m}}$ might not be well-defined for tidally stripped halos, the recovered $M_{\mathrm{200m}}$ values are compared to the peak virial mass achieved throughout the halos’ evolutionary history,
@@ 4.1.2 Mass and structural parameters
    The virial mass: $M_{\mathrm{200m}}\equiv M(r_{\mathrm{200m}})$, where $r_{\mathrm{200m}}$ is the radius at which the mean halo density is 200 times the mean matter density of the Universe.
@@
Simple Jeans $0.11\pm 0.01$ $0.02\pm 0.00$ $-0.46\pm 0.09$
GraphNPE $0.11\pm 0.01$ $0.02\pm 0.00$ $-0.59\pm 0.08$
$\log_{10}(M_{\mathrm{200m}}$ / $\mathrm{M_{\odot}}$)
@@
To calculate the predicted values of $V_{\mathrm{max}}$ and $M_{\mathrm{200m}}$, we extrapolate the DM density profiles to $r_{\mathrm{max}}$ and $r_{\mathrm{200m}}$, respectively. Since many Milky Way dwarf galaxies, as well as those in our test samples, have likely experienced some degree of tidal stripping, the outer regions of their DM halos have already been significantly stripped, and the present-day $M_{\mathrm{200m}}$ is not well-defined. Therefore, we compare our mass estimates to the peak virial mass of the halos throughout their evolutionary history, $M_{\mathrm{200m}}^{\mathrm{peak}}$.
@@
To further quantify the performance of each method, we compute standard performance metrics, including the absolute error (AE), squared error (SE), and negative log-likelihood (NLL). Table 3 presents the mean and the standard error for each metric across all selected galaxies. The AE and SE metrics are computed between the true values and predicted values (which we take to be the median of the posteriors) with the SE metric more sensitive to outliers. The NLL metric, on the other hand, takes into account the overall shape of the posterior. If the posterior were perfectly Gaussian, the NLL would be reduced to a chi-square statistic, with an additional normalization term that accounts for the standard deviation. However, since the posteriors of $V_{\mathrm{max}}$ and $M_{\mathrm{200m}}$ exhibit long tails, as hinted from the asymmetric error bars in Figure 3, the NLL provides a more robust assessment. To compute the NLL, we first fit the posterior using Gaussian Kernel Density Estimation [^9] We use a KDE bandwidth of 1 for $V_{\mathrm{max}}$ and 0.1 for $M_{r_{\star,1/2}}$ and $M_{\mathrm{200m}}$. (KDE) and then evaluate it at the true value.
Both GraphNPE and Jeans modeling recover $V_{\mathrm{max}}$ and $M_{\mathrm{200m}}$ within the 95% confidence intervals. However, it is evident from Figure 3 and Table 3 that GraphNPE provides more accurate and tighter constraints compared to Jeans modeling. Specifically, Jeans tends to overestimate $V_{\mathrm{max}}$ and $M_{\mathrm{200m}}^{\mathrm{peak}}$, while also producing significantly wider 68% confidence intervals. As shown in Table 3, Jeans performs worse according to all three performance metrics: AE, SE, and NLL.
@@ 4.2 Higher-moment methods
This reparameterization ensures that $\tilde{\beta}$ remains finite, as it is constrained to $[-1,1]$, whereas $\beta$ can diverge to $\infty$- ∞. The priors of GravSphere are uniform over $\tilde{\beta}_{0}\in[-0.5,1]$, $\tilde{\beta}_{\infty}\in[-0.5,1]$, $\eta\in[0,3]$, with a fixed $r_{a}=1\,\mathrm{kpc}$. For comparison, the priors of GraphNPE in Table 2 roughly translate to a range of $\tilde{\beta}_{0}\in[-0.2,1]$ and $r_{a}\in[2\times 10^{-3},10^{3}]\,\mathrm{kpc}$, with $\tilde{\beta}_{\infty}=1$ and $\eta=2$ fixed by the definition of the OM profile.
Additionally, GravSphere uses the coreNFWtides model for the DM density profile (Collins et al., 2021) and a “three Plummer” model for the tracer density profile (Rojas-Niño et al., 2016; Read & Steger, 2017). For a more detailed description of the model, we refer readers to Section 4.1 in Collins et al. (2021). Here, we summarize the asymptotic behavior of the coreNFWtides profile. The DM density follows $\rho_{\mathrm{dm}}^{\mathrm{cNFWt}}\propto r^{-\gamma}$, where $\gamma$ transitions from $\gamma_{\mathrm{in}}$ at small radii to $\gamma_{\mathrm{out}}$ at large radii. Both the inner and outer slopes, $\gamma_{\mathrm{in}}$ and $\gamma_{\mathrm{out}}$, are free parameters, with prior distributions $[0,1]$ and $[3.01,5]$ respectively. [^11] The parameters $\gamma_{\mathrm{in}}$ and $\gamma_{\mathrm{out}}$ are equivalent to $n$ and $\delta$ in $\rho_{\mathrm{dm}}^{\mathrm{cNFWt}}$. We use this notation for consistency with the gNFW profile. In comparison, the gNFW profile in GraphNPE follows $\rho_{\mathrm{dm}}^{\mathrm{gNFW}}\propto r^{-\gamma}$ at small radii, where $\gamma$ can vary between $[-1,2]$ (see Table 2), and always transition to $\rho_{\mathrm{dm}}^{\mathrm{gNFW}}\propto r^{-3}$ at large radii.

### v2

@@ 2.2 Test samples of dwarf galaxies
Given that our goal is to apply GraphNPE on Milky Way satellite galaxies, we will focus our studies on the satellites of the Milky Way-mass simulated galaxies. In FIRE-2, halos and subhalos are identified using a modified version of the rockstar  [^5] https://bitbucket.org/awetzel/rockstar-galaxies six-dimensional halo finder (Behroozi et al., 2013), which accounts for multi-mass and multi-species particles. For each subhalo, we use the DM particles assigned by rockstar and identify its associated star particles. We use a similar procedure to Necib et al. (2019) to assign member star particles: we require that the star particles lie within the current $r_{\mathrm{200m}}$ (the radius within which the average density is 200 times the mean matter density of the Universe) of the subhalo, and have velocities within $3\sigma$ of the subhalo’s stellar velocity dispersion.
@@
In addition, we select subhalos with halo-to-stellar mass ratios satisfying $M_{\mathrm{halo}}/M_{\mathrm{\star}}>10^{2}$ and with the number of star particles between $N_{\mathrm{\star}}\in[20,5000]$. This roughly corresponds to a stellar mass range of approximately $M_{\mathrm{\star}}\in[10^{5},10^{7}]\,\mathrm{M_{\odot}}$ and a halo mass range of $M_{\mathrm{halo}}\in[10^{7},10^{10}]\,\mathrm{M_{\odot}}$. [^6] Peak halo mass range of $M^{\mathrm{peak}}_{\mathrm{halo}}\in[10^{8},10^{10}]\,\mathrm{M_{\odot}}$. The number of stars is chosen to encompass observations of classical and faint dwarfs. For instance, the highest number of resolved stars observed for Milky Way dwarfs, Fornax and Sculptor, are 2483 and 1365 stars, respectively (Walker et al., 2009a).
Next, we center each subhalo and compute the positions and velocities of star and DM particles of each subhalo relative to the subhalo’s DM barycenter. We identify the center of each subhalo using the shrinking sphere method (Power et al., 2003) on the DM particles only. Starting with an initial guess for the center, we iteratively compute the DM barycenter of all particles within a specified radius. At each step, the radius is reduced by 10%, and the center is updated to the newly calculated barycenter. The iterations proceed until the center converges to within 0.1% of the previous iteration. In rare cases, when a subhalo is located too close to its host or another subhalo, the computed center may still be offset. These cases are manually checked.
Additionally, the stellar and DM centers are not always perfectly aligned, potentially complicating the analysis. To quantify this misalignment, we compute the position and velocity offsets between the stellar and DM centers for all subhalos in our sample. The median position and velocity offsets are $0.17\,\mathrm{kpc}$ and $1.77\,\mathrm{km/s}$, respectively, with maximum values of $0.66\,\mathrm{kpc}$ and $9.52\,\mathrm{km/s}$. [^7] We note one extreme outlier with position and velocity offsets of $13.7\,\mathrm{kpc}$ and $21.5\,\mathrm{km/s}$, respectively, which was excluded from these statistics. For simplicity, we assume these centers coincide.
@@ 3.1 Training Simulation & Forward Model
We assume the galaxies to be spherically symmetric and in dynamical equilibrium (i.e., the distribution function $f(\vec{x},\vec{v};\theta)$ does not explicitly depend on time). These are common assumptions in traditional Jeans analysis, but as previously discussed, might not be robust against realistic Milky Way dwarfs. Ideally, one can construct a more realistic family of distribution functions from hydrodynamic simulations, such as FIRE-2. However, such simulations are computationally demanding and currently impractical for generating the large training sets required for our framework. Moreover, as we will demonstrate in Section 7.1, even high-resolution zoom-in simulations struggle to capture the observed population of dwarf galaxies, particularly the ultra-faints.
@@
where $\sigma_{r}$ and $\sigma_{t}$ are the radial and tangential velocity dispersions. The velocity anisotropy ranges from -$\infty$ to $1$, where $\beta=1,0,$-$\infty$ corresponds to a radial, isotropic, and tangential velocity profile, respectively. We assume a functional form for $\beta(r)$ following the Osipkov-Merritt (OM) profile (Osipkov, 1979; Merritt, 1985),
@@
where $\beta_{0}$ is the normalization and $r_{a}$ is the anisotropy scale radius. The scale radius $r_{a}$ thus determines the transition from $\beta_{0}$ at small radii to a radially-biased orbits at larger radii. In contrast to the N23 model, which always set $\beta_{0}$ to 0, we let $\beta_{0}$ be a free parameter within the range $\left[-0.5,1\right]$.
@@
$\log_{10}(r_{\mathrm{dm}}/\mathrm{kpc})$ DM scale radius gNFW (Eq. 2) [2,2][-2,2]
$\gamma$ DM inner slope gNFW (Eq. 2) [1,2][-1,2]
@@
$\beta_{0}$ Velocity anisotropy normalization OM (Eq. 6) [0.5,1][-0.5,1]
@@
In total, the current model has three DM parameters ($\rho_{0}$, $r_{s}$, $\gamma$) and three stellar parameters ($\beta_{0}$, $_{a}, rr_{\star}$). We expand the prior range for these parameters from N23 and summarize them in Table 2. Additionally, we select only galaxies with 3-D velocity dispersions within $0.1-50\,\mathrm{km/s}$ for the final training set, which should sufficiently include the population of classical and ultra-faint dwarfs (Pace, 2024).888 Assuming an isotropic velocity distribution, the 3-D velocity dispersion $\sigma_{\mathrm{3-D}}$ is related to the line-of-sight velocity dispersion $\sigma_{\mathrm{los}}$ via $\braket{\sigma_{\mathrm{3-D}}^{2}}=3\braket{\sigma_{\mathrm{los}}^{2}}$. For $\sigma_{\mathrm{3-D}}$ in the range of $(0.1-50)\,\mathrm{km/s}$, $\sigma_{\mathrm{los}}$ is approximately within $(0.06-28.9)\,\mathrm{km/s}$. Among Milky Way dwarfs, the highest and lowest observed line-of-sight velocity dispersions–excluding the Small and Large Magellanic Clouds (SMC and LMC)–is $12.1\pm 0.2\,\mathrm{km/s}$ for Fornax (Walker et al., 2009a; Górski et al., 2011; Muñoz et al., 2018; Wang et al., 2019) and $1.2^{+0.9}_{-0.6}\,\mathrm{km/s}$ for Tucana V (Hansen et al., 2024).
@@ 4.1.1 DM density and velocity anisotropy profiles
Figure 2: Comparison between the inferred velocity anisotropy profiles $\beta(r)=1-\sigma^{2}_{t}/\sigma^{2}_{r}$, where $\sigma_{t}$ and $\sigma_{r}$ are the tangential and radial velocity dispersions, from GraphNPE (orange) and Jeans modeling (blue) for a selection of FIRE-2 galaxies. Values of $\beta=0,1$-$\infty$ correspond to isotropic, radially-biased, and tangentially-biased orbits. Panels are the same as in Figure 1.
Figure 1 and Figure 2 show the recovered DM density profiles $\rho_{\mathrm{dm}}(r)$ and velocity anisotropy profiles $\beta(r)$ for a subset of the selected galaxies, with the remaining profiles provided in Appendix C.1. Each panel shows a single dwarf galaxy, labeled with its rockstar halo ID at $z=0$ and a corresponding letter (e.g., A, B, C, D) for readability, and compares the profiles derived using GraphNPE (orange) and Jeans modeling (blue). The posterior median profiles are shown as solid lines, with shaded bands indicating the 68% and 95% confidence intervals, while black circles with error bars represent the true profiles. [^9] Note that error bars are not visible for the true density profiles. The dashed vertical lines indicate $(0.25,1,4)\,r_{\mathrm{1/2}}$, where $r_{\mathrm{1/2}}$ is the 3-D half-stellar mass radius.
@@
Figure 3: Comparison between the inferred mass and structural parameters from GraphNPE (orange) and Jeans modeling (blue). From left to right, the columns show the peak circular velocity $V_{\mathrm{max}}$, the half-stellar radius mass $M_{r_{\star,1/2}}$, and the virial mass $M_{\mathrm{200m}}^{\mathrm{peak}}$. $V_{\mathrm{max}}$ and $M_{\mathrm{200m}}^{\mathrm{peak}}$ are calculated by extrapolated the DM density profiles to $r_{\mathrm{max}}$ and $r_{\mathrm{200m}}$, respectively. The top and bottom rows show the median and fractional residual of the predicted values against the true values, respectively, with the error bars denoting the 64th percentile ranges. The black dashed line represents the one-to-one correlation. Since the virial mass $M_{\mathrm{200m}}^{\mathrm{peak}}$ might not be well-defined for tidally stripped halos, the recovered $M_{\mathrm{200m}}^{\mathrm{peak}}$ values are compared to the peak virial mass achieved throughout the halos’ evolutionary history,
Figure 4: The fractional residual of the predicted mass and structural parameters as a function of the tracer counts $N_{\star}$. From left to right, the panels show the peak circular velocity $V_{\mathrm{max}}$, the half-stellar radius mass $M_{r_{\star,1/2}}$, and the virial mass $M_{\mathrm{200m}}^{\mathrm{peak}}$. The GraphNPE and Simple Jeans results are shown as orange and blue data points, respectively.
@@ 4.1.2 Mass and structural parameters
    The peak virial mass: The virial mass is defined as $M_{\mathrm{200m}}\equiv M(r_{\mathrm{200m}})$, where $r_{\mathrm{200m}}$ is the radius at which the mean halo density is 200 times the mean matter density of the Universe. However, since many Milky Way dwarf galaxies, as well as those in our test samples, have likely experienced some degree of tidal stripping, the outer regions of their DM halos have already been significantly stripped, and the present-day $M_{\mathrm{200m}}$ is not well-defined. Therefore, we compare our mass estimates to the peak virial mass of the halos throughout their evolutionary history, $M_{\mathrm{200m}}^{\mathrm{peak}}$. Unless stated otherwise, the virial mass will refer to this peak value throughout our analysis.
Note that although the inner slope of the DM density profiles are of considerable interest, we do not include it in our analysis for the following reasons. Since the simulated profiles do not strictly follow gNFW, we estimated the slope using the innermost radial bins. However, this approach yielded noisy and unreliable estimates due to several factors: (1) strong dependence on the binning scheme, (2) limited numbers of DM particles in the innermost regions, particularly for lower-mass subhalos, introducing significant Poisson noise, and (3) finite resolution effects that prevent probing sufficiently deep into the inner profiles. Given these limitations, we concluded that reliable constraints on the inner slope would require either higher-resolution simulations or alternative fitting approaches.
@@
Simple Jeans $11\pm 0.01 0.0$$02\pm 0.00$ $0.46\pm 0.09$
GraphNPE $11\pm 0.01 0.0$$02\pm 0.00$ $0.59\pm 0.08$
$\log_{10}(M_{\mathrm{200m}}^{\mathrm{peak}}$ / $\mathrm{M_{\odot}}$)
@@
Since the model outputs the parameterized gNFW density profiles in Equation 2, for each profile, we calculate the predicted peak circular velocity, half-stellar radius total mass, and peak virial mass by integrating the profiles to their respective radii, i.e. $r_{\mathrm{max}}$, $r_{\star,1/2}$, $r_{\mathrm{200m}}$.
@@
To further quantify the performance of each method, we compute standard performance metrics, including the absolute error (AE), squared error (SE), and negative log-likelihood (NLL). Table 3 presents the mean and the standard error for each metric across all selected galaxies. The AE and SE metrics are computed between the true values and predicted values (which we take to be the median of the posteriors) with the SE metric more sensitive to outliers. The NLL metric, on the other hand, takes into account the overall shape of the posterior. If the posterior were perfectly Gaussian, the NLL would be reduced to a chi-square statistic, with an additional normalization term that accounts for the standard deviation. However, since the posteriors of $V_{\mathrm{max}}$ and $M_{\mathrm{200m}}^{\mathrm{peak}}$ exhibit long tails, as hinted from the asymmetric error bars in Figure 3, the NLL provides a more robust assessment. To compute the NLL, we first fit the posterior using Gaussian Kernel Density Estimation [^10] We use a KDE bandwidth of 1 for $V_{\mathrm{max}}$ and 0.1 for $M_{r_{\star,1/2}}$ and $M_{\mathrm{200m}}$. (KDE) and then evaluate it at the true value.
Both GraphNPE and Jeans modeling recover $V_{\mathrm{max}}$ and $M_{\mathrm{200m}}^{\mathrm{peak}}$ within the 95% confidence intervals. However, it is evident from Figure 3 and Table 3 that GraphNPE provides more accurate and tighter constraints compared to Jeans modeling. Specifically, Jeans tends to overestimate $V_{\mathrm{max}}$ and $M_{\mathrm{200m}}^{\mathrm{peak}}$, while also producing significantly wider 68% confidence intervals. As shown in Table 3, Jeans performs worse according to all three performance metrics: AE, SE, and NLL.
@@
Lastly, we examine the fractional residuals of the mass and structural parameters as a function of the tracer counts $N_{\star}$ in Figure 4, comparing GraphNPE and Simple Jeans results. As expected, GraphNPE demonstrates consistent performance even at low tracer counts ($N_{\star}<100$), while Simple Jeans shows degraded performance in this regime. This is especially true for $V_{\mathrm{max}}$ and $M_{\mathrm{200m}}^{\mathrm{peak}}$ recovery. At high tracer counts ($N_{\star}>100$), the residuals become more similar, though GraphNPE still slightly outperforms Simple Jeans. We present additional results for GraphNPE across the full test dataset in Appendix C.2.
@@ 4.2 Higher-moment methods
Figure 8: The fractional residual of the predicted mass and structural parameters as a function of the tracer counts $N_{\star}$. From left to right, the panels show the peak circular velocity $V_{\mathrm{max}}$, the half-stellar radius mass $M_{r_{\star,1/2}}$, and the virial mass $M_{\mathrm{200m}}^{\mathrm{peak}}$. The GraphNPE and GravSphere results are shown as orange and blue data points, respectively.
@@
This reparameterization ensures that $\tilde{\beta}$ remains finite, as it is constrained to [1,1][-1,1], whereas $\beta$ can diverge to -$\infty$. The priors of GravSphere are uniform over $\tilde{\beta}_{0}\in[-0.5,1]$, $\tilde{\beta}_{\infty}\in[-0.5,1]$, $\eta\in[0,3]$, with a fixed $r_{a}=1\,\mathrm{kpc}$. For comparison, the priors of GraphNPE in Table 2 roughly translate to a range of $\tilde{\beta}_{0}\in[-0.2,1]$ and $r_{a}\in[2\times 10^{-3},10^{3}]\,\mathrm{kpc}$, with $\tilde{\beta}_{\infty}=1$ and $\eta=2$ fixed by the definition of the OM profile.
Additionally, GravSphere uses the coreNFWtides model for the DM density profile (Collins et al., 2021) and a “three Plummer” model for the tracer density profile (Rojas-Niño et al., 2016; Read & Steger, 2017). For a more detailed description of the model, we refer readers to Section 4.1 in Collins et al. (2021). Here, we summarize the asymptotic behavior of the coreNFWtides profile. The DM density follows $\rho_{\mathrm{dm}}^{\mathrm{cNFWt}}\propto r^{-\gamma}$, where $\gamma$ transitions from $\gamma_{\mathrm{in}}$ at small radii to $\gamma_{\mathrm{out}}$ at large radii. Both the inner and outer slopes, $\gamma_{\mathrm{in}}$ and $\gamma_{\mathrm{out}}$, are free parameters, with prior distributions $[0,1]$ and $[3.01,5]$ respectively. [^12] The parameters $\gamma_{\mathrm{in}}$ and $\gamma_{\mathrm{out}}$ are equivalent to $n$ and $\delta$ in $\rho_{\mathrm{dm}}^{\mathrm{cNFWt}}$. We use this notation for consistency with the gNFW profile. In comparison, the gNFW profile in GraphNPE follows $\rho_{\mathrm{dm}}^{\mathrm{gNFW}}\propto r^{-\gamma}$ at small radii, where $\gamma$ can vary between [1,2][-1,2] (see Table 2), and always transition to $\rho_{\mathrm{dm}}^{\mathrm{gNFW}}\propto r^{-3}$ at large radii.